OpenStudy (anonymous):

will give a medal and fan!!!!!!!!!! A parking garage downtown offers two different plans for commuters who have to pay for parking. Plan A charges $3.00 for each day of parking. Plan B requires a $20.00 flat fee which covers the commuter's first 10 days of parking, and then $2.50 for each additional day. If a commuter parks 15 days in one month, which is the better plan and how much does it cost? A. Plan B: $37.50 B. Plan B: $32.50 C. Plan A: $30.00 D. Plan A: $45.00

OpenStudy (anonymous):

Plan A:\[C_A=3x=3(15)=30\] Plan B:\[C_B=2.50(x-10)+20=2.5(5)+20=32.5\] Plan A is cheaper at $30. (C)

OpenStudy (anonymous):

thank you!

OpenStudy (anonymous):

im already a fan of you tho